The lower end of a capillary tube of radius 1 mm is dipped vertically into mercury. (a) Find the depression of mercury column in the capillary. (b) If the length dipped inside is half the answer of part (a) , find the angle made by the mercury surface at the end of the capillary with the vertical. Surface tension of mercury = 0.465 N m-1 and the contact angle of mercury with glass = 130°.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. a) h = 2T cos θ/rpg b) T * 2πr cos θ = πr2h * p * g ∴ cos θ = hrpg/2T